OpenStudy (anonymous):

phosphorus has an atomic mass of 31 and an atomic number of 15 ,so the number of nuetrons must be ?

OpenStudy (anonymous):

16.

OpenStudy (anonymous):

The atomic mass is the atomic number (which is the same as the number of protons in teh nucleus), plus the number of neutrons. So if it has an atomic mass of 31, and an atomic number of 15, then it will have 31-15 = 16 neutrons in the nucleus.

OpenStudy (anonymous):

atomic mass = no. of protons +no. of neutrons => no. of neutrons=atomic mass-no. of protons =31-15 =16